Circular No. 5675

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SWIFT-TUTTLE (1992t) L. K. Deutsch, University of California at Berkeley and Ames Re-

search Center, NASA; W. F. Hoffmann, University of Arizona; J. L. Hora, University of Hawaii; J. R. Regester, University of Arizona; G. G. Fazio, Harvard-Smithsonian Center for Astrophysics; and K. Shivanandan, Center for Advanced Space Sensing, communicate: “We report imaging of P/Swift-Tuttle with the UA/SAO/NRL Mid-Infrared Array Camera (MIRAC; scale 0′′.67/pixel) and the Steward 2.3-m telescope at Kitt Peak. Images were obtained between Dec. 9.04 and 9.09 UT. A preliminary analysis gives the following flux levels and peak surface brightnesses in their respective wavelengths and bandpasses for the comet in a circular region of diameter 6′′ centered on the infrared peak: 49.3 Jy, 6.2 Jy arcsec^{−2}, 8.8 μm, 0.87 μm; 55.7, 6.1, 9.8, 0.96; 76.8, 7.4, 11.7, 1.13; 66.6, 6.8, 12.5, 1.16. The images were calibrated using the infrared flux standard α Tau. There is no evidence for significant silicate emission; the four quoted fluxes for the region centered on the peak are consistent with a 345-K blackbody curve. The strongest mid-infrared emission extends to the southwest in a tail from the peak along p.a. 225^{o}. This structure is surrounded by lower-level emission, which encompasses the peak and extends to the south and west for several arcsec and then begins to curve toward the southeast. For comparison, the measured surface brightnesses at a position 6′′ to the southwest of the peak (p.a. 225^{o}) are 1.6 Jy arcsec^{−2} (8.8 μm); 1.8 (9.8); 2.1 (11.7); 1.8 (12.5). Emission to the west, south, and east extended beyond the edges of our images, which correspond to 18′′.7 west, 13′′.9 south, and 20′′.8 east of the peak in our 11.7-μm image. Emission to the north-northeast extended only 3′′-4′′ from the peak at all wavelengths. The minimum detected surface brightness in our 11.7-μm image was 0.14 Jy arcsec^{−2}.”

HR 1469 = NSV 1671 W. Wenzel, Sonneberg Observatory, reports that no further minimum

(cf. IAUC 5656) and no other conspicuous variability could be detected on any of 1045 photographs taken from 1928 to 1992 Nov. 1.0 UT.

PERIODIC COMET SCHWASSMANN-WACHMANN 1 H. Mikuz, Ljubljana, Slovenia, provides the following additional total

V magnitudes (cf. IAUC 5674) obtained with 0.2-m short-focus cameras (+ CCD): Dec. 13.78 UT, 14.1 (2′.5 tail in p.a. 25^{o}); 14.76, 14.3.

1992 December 15 Daniel W. E. Green